OpenStudy (anonymous):

solve the following system of equations using (i) Gauss - Jacobi iterative technique (ii) Gauss - Seidel iterative technique 10x - 5y -2z = 3 4x -10y + 3z =-3 x + 6y + 10z = -3

OpenStudy (anonymous):

any idea for solving this ??

OpenStudy (mathmale):

Hello, Alooy! I'm very familiar with the Gauss-Jordan method, in case you're interested. The two other methods you've mentioned are probably just variations of Gauss-Jordan. Would it help you at all if I led you through the Gauss-Jordan method to solve this system of linear equations?

OpenStudy (anonymous):

yeah plz can u led me through it @mathmale

OpenStudy (mathmale):

Please make up an "augmented matrix" by taking all of the coefficients of this system of linear equations 10x - 5y -2z = 3 4x -10y + 3z =-3 x + 6y + 10z = -3 For example, the first row of this "augmented matrix" would be 10 -5 -2 3. Have you done this before?

OpenStudy (anonymous):

actually i dont remember :(

OpenStudy (mathmale):

What course are you in? Are you in India or in the USA?

OpenStudy (anonymous):

im in malaysia actually i took leave from classes like tow weeks Illness and i just come back yesterday and im lost now

OpenStudy (anonymous):

im doing electrical and electronic of engineering

OpenStudy (mathmale):

OK, Alooy, let's try solving that system of linear equations using the Gauss-Jordan method. Unfortunately, there are a lot of steps involved. For whatever it's worth I'd like to lead you through the solution and then answer any questions about the process.

OpenStudy (anonymous):

okay sure

OpenStudy (mathmale):

Given the system 10x - 5y -2z = 3 4x -10y + 3z =-3 x + 6y + 10z = -3 Please write / draw the "augmented matrix" by taking the 12 coefficients (without the variables x, y and z): 10 -5 -2 3 4 -10 3 -3 Would you please complete this by copying these first two rows and writing in the third row.

OpenStudy (anonymous):

10 -5 -2 3 4 -10 3 -3 1 6 10 -3

OpenStudy (mathmale):

Good. Now imagine you have square brackets enclosing this 3-row, 4-column matrix. Your matrix should look like this (not finished):